A 35.0 g block of copper at 350.0 degrees celcius is added to a beaker containing 525 g of water at 35.0 degrees celcius. What is the final temperature of water?
1
Expert's answer
2012-11-30T06:36:56-0500
Let us denote
mc - mass of copper = 35 g
cc - specific heat capacity of copper = 385 J/(kg*K)
tc - temperature of copper = 350 C = 623 K
mw - mass of water = 525g
cw - specific heat capacity of water = 4200 J/(kg*K)
tw - temperature of water = 35 C = 308 K

T - final temperature.
we can write

mc*cc*(tc-T) = mw*cw*(T-tw)

Hence
T = (mc*tc*cc + mw*tw*cw) / (mw*cw + mc*tc) = 310K = 37 degrees celcius